OpenStudy (anonymous):

PLEASE HELP ! :) the Pleasant Valley Library is moving to a larger building. To get ready for the move the staff is packing all of the books into boxes. The library owns 35000 books. On average, each book measures 8 in by 10 in yb 2in. Each box measures 2 ft by 2.5 ft by 1.5 ft. 1.About how many boxes will it take to store all of the books? Round to the nearest box.

OpenStudy (anonymous):

Volume of book = Volume of 35,000 bks = volume of box = .. Divide volume of 35,000 bks for box => number of boxes!

OpenStudy (anonymous):

Alright! First, we're going to see how big our total bunch of books is. This means, we're going to multiply the volume of one book (length * width * height) by the total number of books. 8in*10in*2in = 160in^3 per book 160in^3 * 35000 books = 5,600,000in^3 for all the books Now, we're going to find the volume of one box, but first we're going to convert the dimensions of the box to inches. 2ft = 24 in, 2.5ft = 30in, 1.5ft = 18in 24in*30in*18in = 11520 in^3 Now, we divide the total volume of the books by the volume of one box, and this will tell us how many boxes we need. 5600000/11520 = 486.11, but since we can't have part of a box, we round up to 487 boxes.

OpenStudy (anonymous):

@echec's instruction so well at 1ft = 12 in!

OpenStudy (anonymous):

Thank you! Remembering to convert is the hardest part for me. :D

OpenStudy (anonymous):

there's another thia truck to carry the boxes 2. the library uses a truck to carry the boxes of books to the new location. The cargo space in the truck measures 1.5 yards by 4 yards by 2.5 yards. How many tripss will it take for the truck to carry all the books to the new building?

OpenStudy (anonymous):

We're going to do conversions again! Yeah! 1.5 yards = 4.5 ft = 54 in 4 yards = 12 ft = 144 in 2.5 yards = 7.5 ft = 90 in 54in*144in*90in = 699840in^3 We're going to divide this number by the volume of a box, and that will tell us how many boxes fit on one trip. 699840/11520 = 60.75, but because you can't fit .75 of a box on the truck, we round down to 60. If 60 boxes fit per trip, we can divide the number of boxes by 60 and that will tell us how many trips we need. 487 boxes/60 boxes = 8.1166 trips, which rounds up to 9 trips because you can not make just "part" of a trip!

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

Glad to help. :)

OpenStudy (anonymous):

@dariamokhnach One post at a time! Even I want to click on Best Answer for echec, but I can't because you haven't opened new post.

OpenStudy (anonymous):

haha.. next question maybe, i hope echec will answer again :)

OpenStudy (anonymous):

Please mark your question as closed when it is solved.